OpenStudy (clamin):

PLEASE HELP!!! MEDAL!! A rectangular block of copper metal weighs 1896g. The dimensions of the block are 8.4 cm by 5.5 cm by 4.6cm . From this data, what is density of copper?

OpenStudy (clamin):

how do you also calculate for volume???

OpenStudy (anonymous):

\[Density=\frac{ mass }{ volume}\]Unit for density is \[\frac{ g }{cm^3 }\] Mass is g Volume is \[cm^3\] To get cm^3 from cm, you multiply \[cm \times cm \times cm\] That mean volume=width x length x height.